OpenStudy (anonymous):

Hello , Since finals are coming , you guys are going to see me a lot here : D anyways I've gone through some questions and I wanted to check if the answers were right , I'll post them IN a second...

OpenStudy (anonymous):

1: \[\int\limits_{0}^{1} \frac{ 4x +2 }{ (x^2+x+1)^3 }\] My solution: I would do this : take a common number 2 so I'll get: \[\frac{ 2x+1 }{ (x^2+x+1)^3 }\] dx so now the derivative of the denominator equals the numerator , but I don't think it would work since we have a cube , what should I do ? --- 2) \[\int\limits_{1}^{infinite} \frac{ 2 }{ 1+4x^2 }\] dx My solution: I would take 1/4 as a common number and get rid of the 2 then I'l get: \[\frac{ 1 }{ 1/4+x^2 }\] which is \[1/2 \tan^{-1}(2x)\] --- 3) the series from k=0 to infinity for : \[\frac{ 2k^3 }{ 3k^3 -2k +1 } - \frac{ 3 }{ 2^k }\] My solution: since the limit for the first series equals 1 then it's divergent and the other one is also divergent since the first one does. --- and this one I didn't really understand what he wants , could you guys help ? question: Evaluate the area enclosed by the curve corresponding to the following parametric equations: x(t) = sin(3t) y(t)=2cos(3t)-1 0<=t<= pi/2

OpenStudy (sidsiddhartha):

take (x^2+x+1)=u so ( 2x+1)dx=du now just substitute

OpenStudy (anonymous):

So it's -4/(x^2 +x +1)^2

OpenStudy (sidsiddhartha):

no \[2\int\limits_{1}^{3}\frac{ du }{ u ^{3}}\]

OpenStudy (sidsiddhartha):

do u get it?

OpenStudy (anonymous):

Ok , if it was du/u^3 we can't use lnu^3 since it is not equal to du " the derivative "

OpenStudy (shamim):

no

OpenStudy (shamim):

(u^-3)du

OpenStudy (sidsiddhartha):

yeah we cant use that

OpenStudy (anonymous):

I'll solve it after finishing the series chapter , hopefully I would get it , could you guys check the other ones ? specially the last one since I didn't get it at all.

OpenStudy (shamim):

after integration we will get (u^-3+1)/(-3+1)

OpenStudy (anonymous):

thanks guys , as I said I'm more worried about the last one since that chapter we finished it only in 1 week so I didn't get it at all :D check the other ones please.

OpenStudy (sidsiddhartha):

the graph will look like that http://www.wolframalpha.com/input/?i=intersection+of+y%3Dsin3t+and+y%3D2cos%283t%29-1+graph

OpenStudy (anonymous):

I prefer not to graph :D what's the algebraic way to do this ? graphing could get hard sometimes so it's not part of the exam.

OpenStudy (anonymous):

Where are you guys :(

OpenStudy (anonymous):

Usually have to create a post for each question since it might be assumed you only asked 1 question and that it was answered, close the question and ask a new question and it will most likely get looked at

OpenStudy (anonymous):

Oh , thanks buddy , I think i'll close this one then :D